Status Report

Hubble Space Telscope Daily Report #3010 – 3 Dec 2001

By SpaceRef Editor
December 3, 2001
Filed under , ,



PERIOD COVERED: 0000Z (UTC) 12/03/01 – 0000Z (UTC) 12/04/01

Daily Status Report as of 338/0000Z


1.1 Completed STIS/CCD/MA1/MA2 9181 (Exploratory Observations of a New
Bright Quasar)

The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was
used to observe a new, extremely bright quasar that is the second brightest
optical object in the sky. Optically classified as a Narrow-line Seyfert 1
galaxy {NLS1}, PHL 1811 is not typical of this class of objects because it
is a very weak X-ray source. This proposal will perform short exploratory
UV spectroscopic observations of this new bright quasar that will reveal
its true nature and permit identifcation of potential applications of
future deeper observations. There were no reported problems.

1.2 Completed Three Sets of WF/PC-2 9244 (POMS Test Proposal: WFII
Parallel Archive Proposal Continuation)

The WF/PC-2 was used to perform a generic target version of the
Archival Pure Parallel program. The program will be used to take parallel
images of random areas of the sky, following the recommendations of the
Parallels Working Group. The observations completed with no anomalous

1.3 Completed Two Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3)

The WF/PC-2 was used obtain three dark frames every day to provide
data for monitoring and characterizing the evolution of hot pixels. No
problems were encountered.

1.4 Completed Three Sets of WF/PC-2 8940 (Cycle 10 Earth Flats)

The WF/PC-2 was used to monitor flatfield stability by obtaining
sequences of Earth streak flats to construct high quality flat fields for
the WF/PC-2 filter set. These flat fields will allow mapping of the OTA
illumination pattern and will be used in conjunction with previous internal
and external flats to generate new pipeline superflats. The proposal
completed without incident.

1.5 Completed Two Sets of STIS/CCD 8901 (Dark Monitor-Part 1)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the darks. There was no anomalous activity.

1.6 Completed STIS/CCD 9055 (Direct Detection of an Extrasolar Planet
in Reflected Light)

The Space Telescope Imaging Spectrograph (CCD) was used to observe
HD 209458B that is the first extrasolar planet known to transit the disk of
its parent star. It is proposed to detect directly the gas giant planet
orbiting HD 209458 in reflected light by observing the secondary eclipse,
i.e. when the planet is occulted by the star. There were no reported

1.7 Completed Four Sets of WF/PC-2 8935 (Cycle 10 Standard Darks)

The WF/PC-2 was used to obtain dark frames every week in order to
provide data for the ongoing calibration of the CCD dark current rate, and
to monitor and characterize the evolution of hot pixels. Over an extended
period these data will also provide a monitor of radiation damage to the
CCDs. The proposal completed with no reported problems.

1.8 Completed WF/PC-2 8939 (Cycle 10 Internal Monitor)

The WF/PC-2 was used to calibrate the internal monitor, to be run
weekly to monitor the health of the cameras. No problems were encountered.

1.9 Completed STIS/CCD 8900 (Performance Monitor)

The Space Telescope Imaging Spectrograph (CCD) was used to measure
the baseline performance and commandability of the CCD subsystem. All
exposures are internals. The observations completed normally.

1.10 Completed Two Sets of WF/PC-2 9118 (Tracing the Cosmic Expansion
to z>1 with Type Ia Supernovae)

The WF/PC-2 was used to observe type Ia supernovae in order to
provide evidence for an accelerating universe. The case for cosmic
acceleration rests almost entirely on the observation that the observed SN
Ia at z~0.5 are 0.25 magnitudes fainter than expected for a
non-accelerating Universe. It is proposed to follow five SN Ia in the
range 0.95 0 cosmology, this experiment is a powerful and
straightforward way to assess the reliability of the SN Ia
measurements. In addition, if SN Ia are reliable standard candles, the
proposed observations will significantly increase the precision with which
Omega_Lambda and Omega_M are measured. There was no reported anomalous

1.11 Completed Four Sets of STIS/CCD 9285 (POMS Test Proposal: STIS
Non-scripted Parallel Proposal Continuation III)

The Space Telescope Imaging Spectrograph (CCD) was used to make low
galactic latitude, non-scripted parallel observations as part of a POMS
test proposal. The observations were completed as planned, and no
anomalies were reported.

1.12 Completed Three Sets of STIS/CCD 8903 (Bias Monitor – Part 1)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at
gain = 4 to build up high-S/N superbiases and track the evolution of hot
columns. There were no problems.

1.13 Completed STIS/CCD 8611 (Observations of Nearby Type Ia Supernovae)

The Space Telescope Imaging Spectrograph (CCD) was used to obtain
UV spectra of five nearby {0.02 < z < 0.08} SNe Ia in the Hubble Flow. The spectra will be taken at weekly intervals over a range in time starting slightly before maximum light and extending to +30 days. These observations will accomplish the following three goals: {1} calibration of the rest frame UV light curves of SNe Ia and an assessment of their potential use as distance indicators through UV light curve shape analyses. {2} improvement in our understanding of the physics of SNe Ia, metallicity/evolutionary effects and correlations between peak brightness and UV spectral features. {3} calibration of the SNe Ia previously observed by HST at high-redshift. This data is crucial for proper cross-filter k-corrections and calibration of the supernova photometry. The observations completed with no reported anomalous activity.


2.1 Guide Star Acquisitions:

Scheduled Acquisitions: 5

Successful: 5

Scheduled Re-acquisitions: 8

Successful: 8

2.2 FHST Updates:

Scheduled: 18

Successful: 18

2.3 Operations Notes:

The CCS engineering status buffer (ESB) limits were adjusted at
337/1111Z. Then, the ESB buffer was dumped and cleared at 337/1345Z. ROP
DF-18A was used for both exercises.

Using ROP SR-1A, the SSR EDAC error counter was cleared at
337/1348Z and at 338/0247Z.

The ephemeris table was uplimked at 338/0022Z per ROP DF-07A.


Continuation of normal science observations and calibrations.

SpaceRef staff editor.